# How Far Do Time Series Foundation Models Paint the Landscape of Real-World Benchmarks ?

### v1

@@ A.1 Models
Model Params (M) Release Date Architecture Reference amazon/chronos-bolt-tiny $\sim$7M 2024 Encoder-Decoder [17] amazon/chronos-bolt-mini $\sim$21M 2024 Encoder-Decoder [17] amazon/chronos-bolt-small $\sim$48M 2024 Encoder-Decoder [17] amazon/chronos-bolt-base $\sim$205M 2024 Encoder-Decoder [17] amazon/chronos-t5-tiny $\sim$8M 2024 Encoder-Decoder [17] amazon/chronos-t5-mini $\sim$20M 2024 Encoder-Decoder [17] amazon/chronos-t5-small $\sim$46M 2024 Encoder-Decoder [17] amazon/chronos-t5-base $\sim$201M 2024 Encoder-Decoder [17] amazon/chronos-t5-large $\sim$709M 2024 Encoder-Decoder [17] google/timesfm-2.0-500m $\sim$500M 2025 Decoder-only [36] LinearRegression classical Linear Model (Baseline)
@@
A.2 Optical Flow Experimental Details
@@
A.3 Time Series Example
@@
A.4 Principal Component Analysis (PCA)
@@
Figure 3: PCA projection for demonstrating the non-similarity between two datasets
As illustrated in Fig. 3, the two time series datasets demonstrate notable differences in their low-dimensional projections: REAL-V-TSFM is distributed more uniformly, whereas M4 displays a clear skew, with its distribution being sparse in the lower-left and upper-right regions of the PCA plane. This observation not only underscores the limited diversity of existing time series datasets, but also demonstrates the feasibility and significance of extracting diverse and information-rich time series directly from video data. Given the massive availability of video sources, the potential for extracting varied and valuable time series data is both substantial and promising.
A.5 Does model size really matter?
@@
Datasets Model MAPE sMAPE Agg. Relative WQL Agg. Relative MASE M4 amazon/chronos-bolt-tiny 5.07 5.14 0.95 0.63 amazon/chronos-bolt-mini 5.01 5.08 0.97 0.62 amazon/chronos-bolt-small 5.00 5.09 0.96 0.62 amazon/chronos-bolt-base 4.93 5.03 1.00 0.63 amazon/chronos-t5-tiny 7.43 7.44 0.59 1.00 amazon/chronos-t5-mini 7.13 7.09 1.48 0.93 amazon/chronos-t5-small 7.26 7.28 1.53 0.97 amazon/chronos-t5-base 7.31 7.34 1.55 0.98 amazon/chronos-t5-large 7.11 7.18 1.56 0.98 google/timesfm-2.0-500m-pytorch 1.90 2.03 0.39 0.23 LinearRegression 14.14 14.60 1.00 1.00 REAL-V-TSFM amazon/chronos-bolt-tiny 7.68 6.66 0.92 0.67 amazon/chronos-bolt-mini 7.95 6.65 0.92 0.67 amazon/chronos-bolt-small 7.50 6.55 0.92 0.67 amazon/chronos-bolt-base 7.94 6.89 0.92 0.67 amazon/chronos-t5-tiny 10.85 9.72 1.70 1.13 amazon/chronos-t5-mini 9.84 9.04 1.47 0.99 amazon/chronos-t5-small 10.39 9.13 1.50 1.01 amazon/chronos-t5-base 9.89 8.76 1.41 0.95 amazon/chronos-t5-large 9.76 8.81 1.42 0.96 google/timesfm-2.0-500m-pytorch 7.32 6.56 0.87 0.63 LinearRegression 17.65 15.01 1.00 1.00
@@
A.6 Objects in REAL-V-TSFM
@@
Figure 4 illustrates the performance of 3 models on videos containing different objects within the REAL-V-TSFM dataset, where the vertical axis represents the normalized sMAPE. It is evident that the performance varies considerably between different objects, which correspond to diverse types of time series. Interestingly, videos involving animals yield motion-derived time series that are particularly difficult to forecast, resulting in notably lower predictive accuracy. In contrast, videos of inanimate and static objects, such as books, tend to produce more predictable time series, leading to higher performance. Overall, the dataset exhibits substantial diversity, which not only facilitates the evaluation of model generalization, but also provides a novel and valuable perspective on extracting time series from large-scale real-world data.

### v2

@@ Appendix A Models
Model Params (M) Release Date Architecture Reference amazon/chronos-bolt-tiny $\sim$7M 2024 Encoder-Decoder (ansari2024chronos) amazon/chronos-bolt-mini $\sim$21M 2024 Encoder-Decoder (ansari2024chronos) amazon/chronos-bolt-small $\sim$48M 2024 Encoder-Decoder (ansari2024chronos) amazon/chronos-bolt-base $\sim$205M 2024 Encoder-Decoder (ansari2024chronos) amazon/chronos-t5-tiny $\sim$8M 2024 Encoder-Decoder (ansari2024chronos) amazon/chronos-t5-mini $\sim$20M 2024 Encoder-Decoder (ansari2024chronos) amazon/chronos-t5-small $\sim$46M 2024 Encoder-Decoder (ansari2024chronos) amazon/chronos-t5-base $\sim$201M 2024 Encoder-Decoder (ansari2024chronos) amazon/chronos-t5-large $\sim$709M 2024 Encoder-Decoder (ansari2024chronos) google/timesfm-2.0-500m $\sim$500M 2025 Decoder-only (das2024decoder) LinearRegression classical Linear Model (Baseline)
@@
Appendix B Optical Flow Experimental Details
@@
Appendix C Time Series Example
@@
Appendix D Principal Component Analysis (PCA)
@@
We align the two time series datasets to the same sequence length and performing min–max normalization on each sequence individually, and then we combined them and applied PCA. Two-dimensional kernel density estimates were then computed separately for the projections of each dataset. By overlaying semi-transparent heatmaps with contour lines in blue (M4) and orange (REAL‑V‑TSFM), we visually depict the distribution shapes, concentration regions, and overlap & divergence patterns of the two datasets in the first two principal component dimensions. As illustrated in Fig. 2, the two time series datasets demonstrate notable differences in their low-dimensional projections: REAL-V-TSFM is distributed more uniformly, whereas M4 displays a clear skew, with its distribution being sparse in the lower-left and upper-right regions of the PCA plane. This observation not only underscores the limited diversity of existing time series datasets, but also demonstrates the feasibility and significance of extracting diverse and information-rich time series directly from video data. Given the massive availability of video sources, the potential for extracting varied and valuable time series data is both substantial and promising.
Appendix E Does model size really matter?
@@
Model Datasets MAPE sMAPE Agg. Relative WQL Agg. Relative MASE amazon/chronos-bolt-tiny REAL-V-TSFM 7.43 ± 17.40 6.66 ± 9.92 0.92 ± 0.88 0.67 ± 0.62 M4-Weekly 7.22 ± 5.33 7.12 ± 5.27 0.91 ± 0.96 0.59 ± 0.55 M4-Daily 5.07 ± 1.03 5.14 ± 4.03 0.95 ± 0.77 0.63 ± 0.59 electricity_D 6.79 ± 3.43 6.65 ± 3.18 0.82 ± 0.41 0.64 ± 0.38 LOOP_SEATTLE_D 6.03 ± 2.39 6.15 ± 2.49 0.82 ± 0.09 0.81 ± 0.09 amazon/chronos-bolt-mini REAL-V-TSFM 7.37 ± 17.22 6.65 ± 10.16 0.92 ± 0.90 0.67 ± 0.63 M4-Weekly 6.88 ± 5.01 6.76 ± 4.90 0.90 ± 0.98 0.57 ± 0.57 M4-Daily 5.01 ± 4.88 5.08 ± 4.78 0.97 ± 0.87 0.62 ± 0.62 electricity_D 6.73 ± 3.26 6.61 ± 3.05 0.83 ± 0.43 0.65 ± 0.39 LOOP_SEATTLE_D 6.69 ± 2.63 6.79 ± 2.73 0.91 ± 0.08 0.90 ± 0.08 amazon/chronos-bolt-small REAL-V-TSFM 7.50 ± 18.80 6.55 ± 9.38 0.92 ± 0.88 0.67 ± 0.62 M4-Weekly 6.64 ± 4.88 6.55 ± 4.86 0.87 ± 0.97 0.55 ± 0.55 M4-Daily 5.00 ± 3.98 5.09 ± 3.88 0.96 ± 0.88 0.62 ± 0.45 electricity_D 6.71 ± 3.16 6.60 ± 2.96 0.81 ± 0.42 0.64 ± 0.38 LOOP_SEATTLE_D 6.58 ± 2.62 6.66 ± 2.70 0.89 ± 0.09 0.88 ± 0.09 amazon/chronos-bolt-base REAL-V-TSFM 7.32 ± 17.06 6.57 ± 9.93 0.93 ± 0.90 0.67 ± 0.63 M4-Weekly 5.72 ± 3.83 5.70 ± 3.83 0.79 ± 0.87 0.50 ± 0.49 M4-Daily 4.93 ± 3.82 5.03 ± 3.22 1.00 ± 0.85 0.63 ± 0.59 electricity_D 6.52 ± 3.03 6.44 ± 2.88 0.80 ± 0.41 0.62 ± 0.35 LOOP_SEATTLE_D 6.65 ± 2.61 6.75 ± 2.70 0.90 ± 0.09 0.89 ± 0.09 amazon/chronos-t5-tiny REAL-V-TSFM 10.40 ± 20.13 9.28 ± 10.85 5.40 ± 30.04 5.37 ± 33.57 M4-Weekly 10.84 ± 8.94 10.41 ± 6.81 1.32 ± 1.30 0.87 ± 0.81 M4-Daily 7.43 ± 3.28 7.44 ± 4.64 0.59 ± 0.89 1.00 ± 1.25 electricity_D 10.12 ± 4.05 9.82 ± 3.82 1.35 ± 0.64 0.97 ± 0.55 LOOP_SEATTLE_D 7.87 ± 3.17 8.09 ± 3.41 1.10 ± 0.25 1.07 ± 0.23 amazon/chronos-t5-mini REAL-V-TSFM 9.36 ± 18.15 8.66 ± 10.26 5.55 ± 32.01 5.46 ± 34.38 M4-Weekly 9.98 ± 6.57 9.77 ± 6.25 1.22 ± 1.13 0.80 ± 0.69 M4-Daily 7.13 ± 4.57 7.09 ± 6.22 1.48 ± 1.32 0.93 ± 0.75 electricity_D 9.69 ± 4.12 9.53 ± 3.77 1.31 ± 0.63 0.93 ± 0.54 LOOP_SEATTLE_D 7.52 ± 3.03 7.74 ± 3.27 1.03 ± 0.20 1.01 ± 0.18 amazon/chronos-t5-small REAL-V-TSFM 9.76 ± 19.56 8.68 ± 10.07 4.89 ± 30.06 4.93 ± 32.30 M4-Weekly 9.96 ± 6.42 9.70 ± 6.12 1.26 ± 1.16 0.81 ± 0.70 M4-Daily 7.26 ± 6.13 7.28 ± 5.22 1.53 ± 1.25 0.97 ± 0.77 electricity_D 9.88 ± 3.74 9.64 ± 3.30 1.34 ± 0.67 0.97 ± 0.58 LOOP_SEATTLE_D 7.54 ± 2.88 7.66 ± 2.98 1.05 ± 0.15 1.02 ± 0.13 amazon/chronos-t5-base REAL-V-TSFM 9.56 ± 21.12 8.38 ± 9.92 5.57 ± 32.11 5.33 ± 33.89 M4-Weekly 9.34 ± 6.41 9.17 ± 6.20 1.24 ± 1.23 0.79 ± 0.75 M4-Daily 7.31 ± 6.28 7.34 ± 6.39 1.55 ± 1.45 0.98 ± 0.78 electricity_D 9.41 ± 3.47 9.21 ± 3.24 1.25 ± 0.63 0.93 ± 0.57 LOOP_SEATTLE_D 7.45 ± 2.86 7.60 ± 2.98 1.03 ± 0.17 1.01 ± 0.15 amazon/chronos-t5-large REAL-V-TSFM 9.32 ± 18.46 8.40 ± 9.69 5.45 ± 31.23 5.58 ± 34.82 M4-Weekly 8.85 ± 6.17 8.70 ± 6.06 1.19 ± 1.20 0.75 ± 0.76 M4-Daily 7.11 ± 7.02 7.18 ± 4.65 1.56 ± 1.26 0.98 ± 0.88 electricity_D 9.18 ± 3.85 8.99 ± 3.44 1.19 ± 0.56 0.88 ± 0.50 LOOP_SEATTLE_D 7.06 ± 2.83 7.18 ± 2.92 0.98 ± 0.16 0.95 ± 0.13 google/timesfm-1.0-200m-pytorch REAL-V-TSFM 58.29 ± 82.96 66.11 ± 70.55 32.20 ± 75.67 32.90 ± 79.88 M4-Weekly 92.12 ± 103.91 73.04 ± 61.86 24.74 ± 61.39 25.38 ± 68.59 M4-Daily 49.19 ± 55.93 55.26 ± 65.22 20.77 ± 75.33 22.55 ± 26.33 electricity_D 36.88 ± 49.90 29.84 ± 33.19 29.84 ± 33.19 4.41 ± 17.42 LOOP_SEATTLE_D 22.93 ± 9.86 19.85 ± 7.61 3.45 ± 1.22 3.09 ± 1.14 google/timesfm-2.0-500m-pytorch REAL-V-TSFM 6.97 ± 16.63 6.24 ± 9.17 0.91 ± 1.02 0.64 ± 0.65 M4-Weekly 8.30 ± 7.08 8.18 ± 6.70 0.85 ± 0.96 0.60 ± 0.62 M4-Daily 1.9 ± 1.08 2.03 ± 2.55 0.39 ± 0.22 0.23 ± 0.25 electricity_D 6.81 ± 3.63 6.77 ± 3.54 0.80 ± 0.44 0.63 ± 0.39 LOOP_SEATTLE_D 6.79 ± 2.77 6.89 ± 2.86 0.92 ± 0.06 0.90 ± 0.07 LinearRegression REAL-V-TSFM 15.52 ± 28.44 14.10 ± 20.21 1.00 ± 0.00 1.00 ± 0.00 M4-Weekly 21.91 ± 23.49 19.96 ± 19.67 1.00 ± 0.00 1.00 ± 0.00 M4-Daily 14.14 ± 24.99 14.6 ± 19.99 1.00 ± 0.00 1.00 ± 0.00 electricity_D 13.56 ± 10.74 13.93 ± 14.34 1.00 ± 0.00 1.00 ± 0.00 LOOP_SEATTLE_D 7.50 ± 2.97 7.63 ± 3.13 1.00 ± 0.00 1.00 ± 0.00
@@
Appendix F Objects in REAL-V-TSFM
@@
Figure 4 illustrates the performance of 3 models on videos containing different objects within the REAL-V-TSFM dataset, where the vertical axis represents the normalized sMAPE. It is evident that the performance varies considerably between different objects, which correspond to diverse types of time series. Interestingly, videos involving animals yield motion-derived time series that are particularly difficult to forecast, resulting in notably lower predictive accuracy. By examining the videos, we observed that animal behaviors often exhibit unpredictable irregularities, particularly in species with a higher degree of freedom in body movement, such as frogs or foxes, which possess four limbs and a movable head. In contrast, species like birds demonstrate more predictable motion patterns during flight, characterized by regular wing flapping, making their motion-derived time series easier for TSFMs to forecast. In contrast, videos of inanimate and static objects, such as books, tend to produce more predictable time series, leading to higher performance. Overall, the dataset exhibits substantial diversity, which not only facilitates the evaluation of model generalization, but also provides a novel and valuable perspective on extracting time series from large-scale real-world data.